OpenStudy (anonymous):

What is (n-5) * 6 when n = 8? n/4

OpenStudy (dtan5457):

Replace that equation's N with 8 and solve normally

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (dtan5457):

sorry (n-5)6

OpenStudy (dtan5457):

not sure where your n/4 goes into play

OpenStudy (anonymous):

(8 -5) 6 and 8/4 8-5 = 3 * 6 = 18 and 8/4 = 2

OpenStudy (dtan5457):

perfect

OpenStudy (dtan5457):

it looks almost as if you didn't even need help lol

OpenStudy (dtan5457):

good job

OpenStudy (anonymous):

Then 18 / 2?

OpenStudy (dtan5457):

im not sure what your question is. n=8 you have 2 serperate questions of 1.(n-5)6 2. n/4 did you just not get your answers?

OpenStudy (anonymous):

no in the question it's set up with a line like this __________ in between

OpenStudy (dtan5457):

\[\frac{ (n-5)6 }{ n/4 }\]

OpenStudy (dtan5457):

like that?

OpenStudy (anonymous):

Yea and I passed my test thanks ^-^

OpenStudy (dtan5457):

yeah, in that case it would just be 18 divided by 2